OpenStudy (mathhelp):

A motorboat can go 8 miles downstream on a river in 20 minutes. To travel the same 8 miles upstream, the boat takes 30 minutes. Find the speed of the current.

OpenStudy (bahrom7893):

I think you have to find speed in both cases and the difference will be the speed of the current: V1 = 8/20 V2 = 8/30 V = |V1-V2|

OpenStudy (mathhelp):

i have an equation set up but can u review it?

OpenStudy (bahrom7893):

sure post it

OpenStudy (mathhelp):

20r+20c=8 and 30r-30c=6

OpenStudy (mathhelp):

=8 not 6

OpenStudy (anonymous):

Do you still need help

OpenStudy (bahrom7893):

so what are u solving for, and i have no idea what those equations mean..

OpenStudy (anonymous):

I think you can turn it into relative velocity problem

OpenStudy (mathhelp):

r= rate c=current

OpenStudy (mathhelp):

solve for currren...question is at the top

OpenStudy (anonymous):

River speed+Boat Speed= Actual Speed

OpenStudy (anonymous):

River speed we denote x for downstream -x for upstream

OpenStudy (mathhelp):

i need it done by the eqution i have set up